# PlatoNeRF: 3D Reconstruction in Plato’s Cave via Single-View Two-Bounce Lidar

### 21 Dec 2023 v1

@@ 1 Introduction
Single-photon lidar systems, implemented with single-photon avalanche diodes (SPADs), offer an opportunity for accurate single-view 3D reconstruction. Lidar systems typically emit light into the scene and measure the time of flight (ToF) of the light to return to the sensor. As illustrated in Figure 1a, this light reflects off the scene multiple times — we refer to each reflection as a “bounce”. While traditional lidar systems only exploit the first bounce of light from the scene back to the sensor, providing accurate absolute depth, recent work has shown that two-bounce time of flight, i.e. the time it takes for light to reflect off the scene two times before returning to the sensor, can enable reconstruction of occluded objects [7]. While promising, a limitation of existing methods is generalization to the lower spatial- and temporal-resolutions of lidars found on consumer devices.
Our method, called PlatoNeRF, addresses the limitations of single-view NeRF with two-bounce lidar and the limitations of two-bounce lidar with NeRF. The goal of PlatoNeRF is to reconstruct visible and occluded geometry from a single view using ToF measurements of two-bounce light from a single-photon lidar. Like Henley et al. [7], we illuminate individual points in the scene with a pulsed laser. Light is reflected off the illuminated points onto the rest of the scene before reflecting to the sensor. This light, referred to as two-bounce light, contains information about both scene depth and the presence of shadows created by the laser. Our experimental setup is described further in Section 3.1. Using the ToF measurements from multiple illumination points, we train NeRF to reconstruct the two-bounce ToF by modeling two-bounce optical paths. The presence or absence of two-bounce light reveals shadows, which allow occluded geometry to be inferred, and its ToF reveals depth. Our method is able to reconstruct 3D geometry with higher accuracy than existing single-view NeRF or lidar methods. Furthermore, using lidar allows our method to operate with higher ambient light and lower scene albedo than RGB methods that exploit shadows. We also demonstrate our method better generalizes to lower spatial- and temporal-resolutions than existing lidar methods due to our use of an implicit representation.
@@
In addition, we prepare a dataset of simulated scenes captured with a single-photon lidar. We use this data to evaluate our method and our baselines. Simulating such data is challenging and requires domain expertise. To lower the barrier to entry for machine learning with single-photon lidars and to drive future research in this direction, we will release this dataset upon publication.
@@ 2 Related Work
Shape from shadows (SfS) provides a physically-accurate way to infer occluded geometry based on the shadows it casts. While traditional methods use shadowgrams, space carving, and probablistic methods to infer SfS [34, 24, 17], in recent years, NeRF has been shown to be an effective representation for learning SfS [40, 20, 46, 19, 14]. These methods leverage volumetric rendering to reconstruct the object or scene based on the observation that pixels in shadow result from geometry between the shadowed point and the light source. However, the performance of these methods degrades when the shadow becomes invisible — either due to ambient light or low albedo backgrounds. In contrast, our method, while still relying on shadows to reconstruct occluded areas, is robust to these effects due to our use of lidar rather than RGB sensors.
@@ 3 NeRF from Single-View Two-Bounce Lidar
In this section, we outline a method to extract 3D geometry from two-bounce transient measurements with NeRF. In Section 3.1, we describe the experimental setup, image formation model, and two-bounce transients. In Section 3.2, we describe how NeRF is trained with supervision from two-bounce transients. In Section 3.3, we provide implementation details that enable replication of our method and results.
@@ 3.1 Notations and Problem Definition
Our experimental setup is shown in Figure 2. The lidar system consists of a SPAD sensor and pulsed laser at known positions $\mathbf{x}_{s}$ and $\mathbf{x}_{l}$ respectively. The laser sequentially points at $K$ different points $\mathcal{A}=\{\mathbf{l}_{1},...,\mathbf{l}_{K}\}$. For each illumination point $\mathbf{l}_{k}$, an image $\mathbf{i}_{k}$ is captured, resulting in a set of $K$ image captures =$\mathcal{I}=\{\mathbf{i}_{1},...,\mathbf{i}_{K}\}$, as illustrated in Figure 1.
@@
SPAD sensors are able to infer properties of a scene by measuring light that has interacted with the scene. In this problem, we are interested in inferring 3D scene geometry from one-bounce and two-bounce light, where “bounce” denotes the number of times light reflects off a scene surface. In Figure 2, light that travels along the path $\mathbf{x}_{l}\rightarrow\mathbf{l}\rightarrow\mathbf{x}_{s}$ is one-bounce light because it undergoes one reflection at $\mathbf{l}$. Similarly, light that travels along the path $\mathbf{x}_{l}\rightarrow\mathbf{l}\rightarrow\mathbf{x}_{p}\rightarrow\mathbf{x}_{s}$ is referred to as two-bounce light because it undergoes two reflections at $\mathbf{l}$ and $\mathbf{x}_{p}$. We refer to each illumination point $\mathbf{l}$ as a virtual source because it acts as a point light source. Similarly, we define $\mathbf{x}_{p}$ as a virtual detector because it refers to the scene point that is observing light from $\mathbf{l}$. In measurement $\mathbf{i}_{k}$, the pixel observing scene point $\mathbf{l}_{k}$ measures one-bounce signal, and all other pixels (e.g. $\mathbf{x}_{p}$) measure two-bounce signals or shadows. In general, one-bounce light arrives at the sensor earlier in time because it travels a shorter optical pathlength. One-bounce light also generally has higher intensity because light intensity is attenuated after every surface reflection.
Figure 3: Method. PlatoNeRF learns 3D scene geometry from single-view two-bounce lidar time of flight, modeled with NeRF. Our method consists of three steps. (a) First, we render primary rays from the camera to the scene (Section 3.2.1). (b) Second, we model rays that scatter and travel to the virtual light (the point where light rays first hit the scene) (Section 3.2.2). Both steps are supervised with transients measured by a single-photon lidar. (c) Third, we find that reconstructing the two-bounce time of flight enables 3D reconstruction (Section 3.2.3).
@@
Each image $\mathbf{i}\in\mathbb{R}^{N_{u}\times N_{v}\times N_{t}}$ is a transient measurement, where $N_{u}$ and $N_{v}$ represent the spatial resolution and $N_{t}$ denotes the number of timing bins. A transient $\mathbf{i}(u,v,t)$ measures the amount of light arriving at every pixel $(u,v)$ at a given time $t$. A pixel measurement $\mathbf{i}(u,v,:)$ measures a histogram of light intensity as a function of time. Each bin of the histogram is discretized to a timing resolution, or bin width, of $t_{\text{res}}$. For our experiments, we choose $t_{\text{res}}=128$ ps, meaning that we can resolve the pathlength of light up to a precision of $3.8$ cm. An example histogram two-bounce signal is plotted in Figure 2. The location of the two-bounce peak $t_{\text{peak}}$ in the histogram is directly correlated to the pathlength $d$ that the light travels via the following equation
@@
where $c\approx 3\cdot 10^{8}$ m/s is the speed of light, $d_{1}$ corresponds to the distance between the laser $\mathbf{x}_{l}$ and the virtual source $\mathbf{l}$, $d_{2}$ is the distance between the virtual source $\mathbf{l}$ and virtual detector $\mathbf{x}_{p}$, and $d_{3}$ is the distance between the virtual detector $\mathbf{x}_{p}$ and sensor $\mathbf{x}_{s}$, as shown in Figure 2.
@@
Equation 2 assumes that a direct path exists between $\mathbf{x}_{p}$ and $\mathbf{l}$. However, if $\mathbf{x}_{p}$ lies in shadow, no two-bounce signal will be measured (i.e., no pulse will be observed). $\mathbf{x}_{p}$ is defined to lie in shadow if an opaque object lies along the ray connecting $\mathbf{l}$ and $\mathbf{x}_{p}$. Because $\mathbf{l}$ is modeled as a point light source, we neglect any diffraction effects and soft shadows that are common with area sources.
@@ 3.2 Two-Bounce Volumetric Lidar Rendering
We parameterize our scene as a neural radiance field (NeRF). The MLP $f_{\theta}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ predicts a volume density $\sigma$ for every input 3D scene point $\mathbf{x}=(x,y,z)$. The 3D geometry of the scene can then directly be estimated from $\sigma(\mathbf{x})$. The goal of this subsection is to synthesize transient images by developing a renderer that can map densities $\sigma$ to predicted transients $\hat{\mathbf{i}}$. These synthesized transient measurements can then be used to train the NeRF in an analysis-by-synthesis framework. Note that, unlike a vanilla NeRF, we are not computing radiance, because we only reconstruct 3D geometry, not texture. Our method is summarized in Figure 3.
@@ 3.2.1 Rendering Primary Rays
The goal of rendering along the primary ray is to compute the two-bounce time-of-flight $t_{\text{peak}}=d/c$ by determining the depth $d_{3}$ of $\mathbf{x}_{p}$. Once the location of $\mathbf{x}_{p}$ is known, $d_{1}$ and $d_{2}$ can subsequently be computed because $\mathbf{l}$, $\mathbf{x}_{s}$, and $\mathbf{x}_{l}$ are already known (Equation 2). First, the MLP is queried at $P$ sampled points along the primary ray $\mathbf{r}_{p}$ between the near plane and far plane to output densities $\sigma_{1},...,\sigma_{P}$. The depth along the ray can be computed from the densities as
@@ 3.2.3 Synthesizing Transient Measurements
Using the synthesized two-bounce time-of-flight and shadow measurements, we can compute a loss based on the input transient measurement. We assume that $\mathbf{l}$, $\mathbf{x}_{s}$, and $\mathbf{x}_{l}$ are known. In addition, we assume that we have a binary mask $\mathbf{m}_{k}\in\mathbb{R}^{N_{u}\times N_{v}}$ for each measurement $\mathbf{i}_{k}$ that segments the transient image into shadowed and unshadowed pixels. Details on how to compute these quantities using the raw transient measurements are further explained in Section 3.3.
@@
The distance loss measures the accuracy of the synthesized two-bounce time of flight. Recall that rendering the primary ray enables estimation of the two-bounce time of flight using Equation 2 and $\mathbf{x}_{p}$ obtained from Equation 3. The distance loss is expressed as
@@
The shadow loss determines if $\mathbf{x}_{p}$ is correctly classified as a shadowed or unshadowed pixel based on the rendered value $p_{\text{shadow}}$. The shadow loss is computed using the output rendering from the secondary ray in Equation 5. The shadow loss is expressed as
@@ 3.3 Implementation Details
Data Pre-Processing. Our method requires five inputs for each pixel: (1) the sensor location $\mathbf{o}_{p}=\mathbf{x}_{s}$ and ray direction $\mathbf{d}_{p}$, (2) laser location $\mathbf{x}_{l}$, (3) the distance from the laser to the virtual source $\|\mathbf{l}-\mathbf{x}_{l}\|$, (4) two-bounce time of flight $t_{\text{peak}}$, and (5) if the pixel is in shadow or not. (1) can be computed using camera matrices and (2) is assumed to be calibrated. We use signal processing to extract (3–5). We compute time-of-flight by using a template of the laser pulse shape as a match filter, and compute the cross-correlation of the match filter with the histogram at every pixel. The peak of the cross-correlation yields the time-of-flight at that pixel, yielding (4), and the maximum value of the cross-correlation yields the confidence that a pulse was measured. We filter out one-bounce returns by finding rays that are parallel or close to parallel to $\mathbf{l}$, and setting the corresponding histogram intensities to zero. Then, (5) can be computed by thresholding the confidence map to yield the binary shadow mask. (3) can be computed by determining the pixel with one-bounce return and computing the distance as $d_{1}=ct_{1B}/2$, where $t_{1B}$ is one-bounce time-of-flight, since the laser and sensor are roughly co-located in our experiments.
@@ 4.1 Datasets
Simulated Datasets. We create datasets of four scenes of a room with either a chair, bunny, dragon, or occluded bunny in a chair, shown in Figure 4. The scenes are captured using a time-of-flight extension of Mitsuba [12] created by Pediredla et al. [29]. For each scene, we heuristically choose $N\!=\!16$ points in the left and right parts of the scene, corresponding to the left and right walls, to illuminate. For each point that is illuminated, we record a transient image. Our scene is measured using a 512$\times$512 SPAD with a temporal resolution of 128 ps (3.84 cm). Intensity is measured for 0.05 $\mu$s (15 m) per illumination spot, resulting in 391 timing bins per pixel. We also render corresponding ground truth depth images both from the training view and across 120 test views around the scene for evaluation. Transient data, depth, and all sensor and illumination parameters will be released.
Real Dataset. We use a dataset of single-photon lidar data captured by Henley et al. [7] to validate our method outside of simulation. The dataset captures a simple indoor scene, shown in Figure 5, containing a mannequin and box. The scene is captured with a 200$\times$200 pixel sensor with an instrument response function of 128 ps (full width at half maximum). The scene is illuminated with 16 laser spots and a per-pixel transient is captured for each laser spot.

### 05 Apr 2024 v2

@@ 1 Introduction
Single-photon lidar systems, implemented with single-photon avalanche diodes (SPADs), offer an opportunity for accurate single-view 3D reconstruction. Lidar systems typically emit light into the scene and measure the time of flight (ToF) of the light to return to the sensor. As illustrated in Fig. 1a, this light reflects off the scene multiple times — we refer to each reflection as a “bounce”. While traditional lidar systems only exploit the first bounce of light from the scene back to the sensor, providing accurate absolute depth, recent work has shown that two-bounce time of flight, i.e. the time it takes for light to reflect off the scene two times before returning to the sensor, can enable reconstruction of occluded objects [7]. While promising, a limitation of existing methods is generalization to the lower spatial- and temporal-resolutions of lidars found on consumer devices.
Our method, called PlatoNeRF, addresses the limitations of single-view NeRF with two-bounce lidar and the limitations of two-bounce lidar with NeRF. The goal of PlatoNeRF is to reconstruct visible and occluded geometry from a single view using ToF measurements of two-bounce light from a single-photon lidar. Like Henley et al. [7], we illuminate individual points in the scene with a pulsed laser. Light is reflected off the illuminated points onto the rest of the scene before reflecting to the sensor. This light, referred to as two-bounce light, contains information about both scene depth and the presence of shadows created by the laser. Our experimental setup is described further in Sec. 3.1. Using the ToF measurements from multiple illumination points, we train NeRF to reconstruct the two-bounce ToF by modeling two-bounce optical paths. The presence or absence of two-bounce light reveals shadows, which allow occluded geometry to be inferred, and its ToF reveals depth. Our method is able to reconstruct 3D geometry with higher accuracy than existing single-view NeRF or lidar methods. Furthermore, using lidar allows our method to operate with higher ambient light and lower scene albedo than RGB methods that exploit shadows. We also demonstrate our method better generalizes to lower spatial- and temporal-resolutions than existing lidar methods due to our use of an implicit representation.
@@
In addition, we prepare a dataset of simulated scenes captured with a single-photon lidar. We use this data to evaluate our method and our baselines. Simulating such data is challenging and requires domain expertise. To lower the barrier to entry for machine learning with single-photon lidars and to drive future research in this direction, we have released this dataset, along with our code and model checkpoints, on our project page.
@@ 2 Related Work
Shape from shadows (SfS) provides a physically-accurate way to infer occluded geometry based on the shadows it casts. While traditional methods use shadowgrams, space carving, and probablistic methods to infer SfS [34, 24, 17], in recent years, NeRF has been shown to be an effective representation for learning SfS [40, 20, 47, 19, 14]. These methods leverage volumetric rendering to reconstruct the object or scene based on the observation that pixels in shadow result from geometry between the shadowed point and the light source. However, the performance of these methods degrades when the shadow becomes invisible — either due to ambient light or low albedo backgrounds. In contrast, our method, while still relying on shadows to reconstruct occluded areas, is more robust to these effects due to our use of lidar rather than RGB sensors.
@@ 3 NeRF from Single-View Two-Bounce Lidar
In this section, we outline a method to extract 3D geometry from two-bounce transient measurements with NeRF. In Sec. 3.1, we describe the experimental setup, image formation model, and two-bounce transients. In Sec. 3.2, we describe how NeRF is trained with supervision from two-bounce transients. In Sec. 3.3, we provide implementation details that enable replication of our method and results.
@@ 3.1 Notations and Problem Definition
Fig. 2 shows our experimental setup. The lidar system consists of a SPAD sensor and pulsed laser at known positions $\mathbf{x}_{s}$ and $\mathbf{x}_{l}$ respectively. The laser sequentially points at $K$ different points $\mathcal{A}=\{\mathbf{l}_{1},...,\mathbf{l}_{K}\}$. For each illumination point $\mathbf{l}_{k}$, an image $\mathbf{i}_{k}$ is captured, resulting in a set of $K$ images =$\mathcal{I}=\{\mathbf{i}_{1},...,\mathbf{i}_{K}\}$, as illustrated in Fig. 1.
@@
SPAD sensors are able to infer properties of a scene by measuring light that has interacted with the scene. In this problem, we are interested in inferring 3D scene geometry from one-bounce and two-bounce light, where “bounce” denotes the number of times light reflects off a scene surface. In Fig. 2, light that travels along the path $\mathbf{x}_{l}\rightarrow\mathbf{l}\rightarrow\mathbf{x}_{s}$ is one-bounce light because it undergoes one reflection at $\mathbf{l}$. Similarly, light that travels along the path $\mathbf{x}_{l}\rightarrow\mathbf{l}\rightarrow\mathbf{x}_{p}\rightarrow\mathbf{x}_{s}$ is referred to as two-bounce light because it undergoes two reflections at $\mathbf{l}$ and $\mathbf{x}_{p}$. We refer to each illumination point $\mathbf{l}$ as a virtual source because it acts as a point light source. Similarly, we define $\mathbf{x}_{p}$ as a virtual detector because it refers to the scene point that is observing light from $\mathbf{l}$. In measurement $\mathbf{i}_{k}$, the pixel observing scene point $\mathbf{l}_{k}$ measures one-bounce signal, and all other pixels (e.g. $\mathbf{x}_{p}$) measure two-bounce signals or shadows. In general, one-bounce light arrives at the sensor earlier in time because it travels a shorter optical pathlength. One-bounce light also generally has higher intensity because light intensity is attenuated after every surface reflection.
Figure 3: Method. PlatoNeRF learns 3D scene geometry from single-view two-bounce lidar time of flight, modeled with NeRF. Our method consists of three steps. (a) First, we render primary rays from the camera to the scene (Sec. 3.2.1). (b) Second, we model rays that scatter and travel to the virtual light (the point where light rays first hit the scene) (Sec. 3.2.2). Both steps are supervised with transients measured by a single-photon lidar. (c) Third, we find that reconstructing the two-bounce time of flight enables 3D reconstruction (Sec. 3.2.3).
@@
Each image $\mathbf{i}\in\mathbb{R}^{N_{u}\times N_{v}\times N_{t}}$ is a transient measurement, where $N_{u}$ and $N_{v}$ represent the spatial resolution and $N_{t}$ denotes the number of timing bins. A transient $\mathbf{i}(u,v,t)$ measures the amount of light arriving at every pixel $(u,v)$ at a given time $t$. A pixel measurement $\mathbf{i}(u,v,:)$ measures a histogram of light intensity as a function of time. Each bin of the histogram is discretized to a timing resolution, or bin width, of $t_{\text{res}}$. For our experiments, we choose $t_{\text{res}}=128$ ps, meaning that we can resolve the pathlength of light up to a precision of $3.8$ cm. An example histogram two-bounce signal is plotted in Fig. 2. The location of the two-bounce peak $t_{\text{peak}}$ in the histogram is directly correlated to the pathlength $d$ that the light travels via the following equation
@@
where $c\approx 3\cdot 10^{8}$ m/s is the speed of light, $d_{1}$ corresponds to the distance between the laser $\mathbf{x}_{l}$ and the virtual source $\mathbf{l}$, $d_{2}$ is the distance between the virtual source $\mathbf{l}$ and virtual detector $\mathbf{x}_{p}$, and $d_{3}$ is the distance between the virtual detector $\mathbf{x}_{p}$ and sensor $\mathbf{x}_{s}$, as shown in Fig. 2.
@@
Eq. 2 assumes that a direct path exists between $\mathbf{x}_{p}$ and $\mathbf{l}$. However, if $\mathbf{x}_{p}$ lies in shadow, no two-bounce signal will be measured (i.e., no pulse will be observed). $\mathbf{x}_{p}$ is defined to lie in shadow if an opaque object lies along the ray connecting $\mathbf{l}$ and $\mathbf{x}_{p}$. Because $\mathbf{l}$ is modeled as a point light source, we neglect any diffraction effects and soft shadows that are common with area sources.
@@ 3.2 Two-Bounce Volumetric Lidar Rendering
We parameterize our scene as a neural radiance field (NeRF). The MLP $f_{\theta}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ predicts a volume density $\sigma$ for every input 3D scene point $\mathbf{x}=(x,y,z)$. The 3D geometry of the scene can then directly be estimated from $\sigma(\mathbf{x})$. The goal of this subsection is to synthesize transient images by developing a renderer that can map densities $\sigma$ to predicted transients $\hat{\mathbf{i}}$. These synthesized transient measurements can then be used to train the NeRF in an analysis-by-synthesis framework. Note that, unlike a vanilla NeRF, we are not computing radiance, because we only reconstruct 3D geometry, not texture. Our method is summarized in Fig. 3.
@@ 3.2.1 Rendering Primary Rays
The goal of rendering along the primary ray is to compute the two-bounce time-of-flight $t_{\text{peak}}=d/c$ by determining the depth $d_{3}$ of $\mathbf{x}_{p}$. Once the location of $\mathbf{x}_{p}$ is known, $d_{1}$ and $d_{2}$ can subsequently be computed because $\mathbf{l}$, $\mathbf{x}_{s}$, and $\mathbf{x}_{l}$ are already known (Eq. 2). First, the MLP is queried at $P$ sampled points along the primary ray $\mathbf{r}_{p}$ between the near plane and far plane to output densities $\sigma_{1},...,\sigma_{P}$. The depth along the ray can be computed from the densities as
@@ 3.2.3 Synthesizing Transient Measurements
Using the synthesized two-bounce time-of-flight and shadow measurements, we can compute a loss based on the input transient measurement. We assume that $\mathbf{l}$, $\mathbf{x}_{s}$, and $\mathbf{x}_{l}$ are known. In addition, we assume that we have a binary mask $\mathbf{m}_{k}\in\mathbb{R}^{N_{u}\times N_{v}}$ for each measurement $\mathbf{i}_{k}$ that segments the transient image into shadowed and unshadowed pixels. Details on how to compute these quantities using the raw transient measurements are further explained in Sec. 3.3.
@@
The distance loss measures the accuracy of the synthesized two-bounce time of flight. Recall that rendering the primary ray enables estimation of the two-bounce time of flight using Eq. 2 and $\mathbf{x}_{p}$ obtained from Eq. 3. The distance loss is expressed as
@@
The shadow loss determines if $\mathbf{x}_{p}$ is correctly classified as a shadowed or unshadowed pixel based on the rendered value $p_{\text{shadow}}$. The shadow loss is computed using the output rendering from the secondary ray in Eq. 5. The shadow loss is expressed as
@@ 3.3 Implementation Details
Data Pre-Processing. Our method requires five inputs per pixel: (1) sensor location $\mathbf{o}_{p}=\mathbf{x}_{s}$ and ray direction $\mathbf{d}_{p}$, (2) laser location $\mathbf{x}_{l}$, (3) distance from the laser to the virtual source $\|\mathbf{l}-\mathbf{x}_{l}\|$, (4) two-bounce time of flight $t_{\text{peak}}$, and (5) if the pixel is in shadow. (1) can be computed using camera matrices and (2) is assumed to be calibrated. We use signal processing to extract (3–5). We compute time-of-flight with a template of the laser pulse shape as a match filter, and compute the cross-correlation of the match filter with the histogram at every pixel. The peak of the cross-correlation yields (4) and the maximum value of the cross-correlation yields the confidence that a pulse was measured. We filter out one-bounce returns by finding rays that are close to parallel to $\mathbf{l}$, and setting the corresponding histogram intensities to zero. Then, (5) can be computed by thresholding the confidence map to yield the binary shadow mask. We use a threshold of $0.15$ in all simulated experiments, except those with ambient light, which require further threshold finetuning. (3) can be computed by determining the pixel with a one-bounce return and computing the distance as $d_{1}=ct_{1B}/2$, where $t_{1B}$ is one-bounce time-of-flight, since the laser and sensor are roughly co-located in our experiments.
@@ 4.1 Datasets
Simulated Datasets. We create datasets of four scenes of a room with either a chair, bunny, dragon, or occluded bunny in a chair, shown in Fig. 4. The scenes are captured using a time-of-flight extension of Mitsuba [12] created by Pediredla et al. [29]. For each scene, we heuristically choose $N\!=\!16$ points in the left and right parts of the scene, corresponding to the left and right walls, to illuminate. For each point that is illuminated, we record a transient image. Our scene is measured using a 512$\times$512 SPAD with a temporal resolution of 128 ps (3.84 cm). Intensity is measured for 0.05 $\mu$s (15 m) per illumination spot, resulting in 391 timing bins per pixel. We also render corresponding ground truth depth images both from the training view and across 120 test views around the scene for evaluation. Transient data, depth, and all sensor and illumination parameters will be released.
Real Dataset. We use a dataset of single-photon lidar data captured by Henley et al. [7] to validate our method outside of simulation. The dataset captures a simple indoor scene, shown in Fig. 5, containing a mannequin and box. The scene is captured with a 200$\times$200 pixel sensor with an instrument response function of 128 ps (full width at half maximum). The scene is illuminated with 16 laser spots and a per-pixel transient is captured for each laser spot.